OpenStudy (anonymous):

Find the distance between the points with coordinates (5, 7) and (-3, 1). (Graph the points on a coordinate plane, if you need to.) 10 100 28 14

OpenStudy (anonymous):

\[ (5, 7) and (-3, 1)\\ \sqrt{(5+3)^2 +(7-1)^2 }=\sqrt{64+36}=\sqrt{100}=10 \]

OpenStudy (anonymous):

We used the distance formula.